Normalize the wavefunction for a 1s atomic orbital of a hydrogen-like atom with atomic number, Z: wEKOxBnJxWfkgAAAABJRU5ErkJggg==, where 6ecp7H7ygQAAAABJRU5ErkJggg==
Answer #1

If the function is normalized then, r yay dł = 1 all space since it is polar coordinate rodisiance from mucleas, has a limit

> N2 93a 1 N = K ra3 -ta (0,) = + / Z317 13e NOTE: According to Gamma function, Tanpa xn-3e-4 dx = (n-1)! Cos (0)= 1 & cos =